OpenStudy (anonymous):

Use the quadratic formula to solve x2 – 4 = 0.

OpenStudy (anonymous):

soo easy use the quadratic formula placed in attachment then put you answer in a,b,c,d order a=1 b=0 c=-4 Then you perform the steps within the quadratic square root which is b^2-4ac \[\sqrt{0^2-4(1)(-4)}= \sqrt{16}= 4\] 4 is the answer

OpenStudy (theviper):

\[x^2-4=0\] =>\[x^2=4\] =>\[x=\sqrt{4}\] =>\[x=2\]

OpenStudy (theviper):

Use the method given by aussy123 @alexxmarie

OpenStudy (anonymous):

oh yea and \[\sqrt{4}=2 \] so sorry 2 is the answer....THANKS VIPER

OpenStudy (theviper):

Anytime:)